# Pluto and Charon: A Time and Memory Efficient Collaborative Edge AI Framework for Personal LLMs Fine-Tuning

### v1

@@ 3. System Overview
PAC is a time, memory and parameter efficient collaborative framework for personal LLMs fine-tuning across multiple edge devices. PAC first equips the target LLM with our Parallel Adapters module (Step 0). PAC profiler fine-tunes the LLM using a calibration dataset on edge devices to record the runtime profile required for parallelism planning (Step 1). PAC planner then takes the profiling results as input and generates planning configurations, including LLM partitioning points and device grouping strategies (Step 2). We configure the Parallel Adapters as trainable while freezing the LLM backbone parameters (Step 3). The parallel configurations generated by the PAC planner are then applied to the edge devices, enabling time, memory, and parameter-efficient hybrid data and pipeline parallelism fine-tuning of the target LLM (Step 4). Since the LLM backbone parameters remain fixed, the intermediate activations generated by the backbone model are invariant for a given input sequence. The PAC maintains a cache of these invariant activations. Through leveraging the cached activations, the efficiency of the fine-tuning process can be accelerated (Step 5).
4. Time, Memory and Parameter Efficient Fine-Tuning Algorithm
4.1. Fine-Tuning LLMs with Parallel Adapters
@@
Observation and Key Insight. As discussed in §2, while techniques such as LoRA (Hu et al., 2021) and Adapters (Houlsby et al., 2019) reduce the number of parameters that need to be updated during fine-tuning, they do not significantly reduce the computational and memory requirements during the training on edge devices. This is because the parameters being updated are still inside the LLM backbone. To calculate the gradients for backpropagation, the full backward passes through the entire pre-trained model are still necessary, as illustrated in Figure 5(a) and (b). In the research field of AI, side-tuning (Zhang et al., 2020) is a specialized fine-tuning technique. It adds a trainable side network that runs in parallel to the backbone model, with the side network’s representation summed with the backbone’s output in the final layer. Crucially, side-tuning only updates the side network, without backpropagating through the backbone model.
@@
Figure 5. Comparison between LLMs fine-tuning with LoRA, Adapters, and our Parallel Adapters.
@@
Parallel Adapters Architecture. In light of side-tuning, we employ a time and memory efficient personal LLMs fine-tuning technique with Parallel Adapters. The overall structure is illustrated in Figure 5(c). Specifically, we decouple conventional Adapters (Houlsby et al., 2019) from the LLM backbone, avoiding their integration at the end of each transformer layer. Instead, we provide a dedicated parallel highway for our trainable adapters network, which takes intermediate activations from the backbone transformer as input and generates the final predictions. In this way, backpropagation through the LLM backbone is free, reducing memory demands for massive activations and computational burdens, thereby enhancing time and memory efficiency over techniques like Adapters and LoRA. Our adapters module demonstrates comprehensive compatibility with established LLM fine-tuning adapters architectures, including the use of linear layers for upward and downward projections as well as trimmed lightweight versions of the backbone transformer (Houlsby et al., 2019; Han et al., 2024; Sung et al., 2022). To ensure the lightweight and resource-efficient nature of our parallel network, the hidden dimension of our Parallel Adapters will be $r$, where $r\ll d$. Specifically, considering an LLM backbone, composed of $L$ layers, and thus $L$ intermediate outputs $\mathrm{b}_{1},\mathrm{b}_{2},\ldots\mathrm{b}_{L}$, each contains $n$ tokens with a hidden dimensionality of $d$, $\mathbf{b}_{i}\in\mathbb{R}^{n\times d}$. We denote the embedding input sequence as $\mathrm{b}_{0}\in\mathbb{R}^{n\times d}$. Assuming adapters are inserted after every layer of backbone, Parallel Adapters consist of $L$ adapters, yielding $L$ intermediate outputs $\mathrm{a}_{1},\mathrm{a}_{2},\ldots\mathrm{a}_{L}$, $\mathbf{a}_{i}\in\mathbb{R}^{n\times r}$. We denote $\mathrm{a}_{0}=\mathbf{W}_{down}\mathrm{b}_{0}$, where $\mathbf{W}_{down}\in\mathbb{R}^{d}\to\mathbb{R}^{r}$. We learn function $f_{i}$ for $i$-th adapter of our Parallel Adapters, which operate on these intermediate outputs.
(1) $\mathrm{a}_{i}=f_{i}(\mathrm{b}_{i},\mathrm{a}_{i-1}).$
Our evaluation in §6 reveals that parallel adapters can achieve comparable model performance to mainstream fine-tuning techniques while being more resource-efficient and better suited for resource-constrained edge environments.
4.2. PAC Activation Cache for Parallel Adapters
@@
Fine-Tuning Parallel Adapters with PAC Activation Cache. Our key idea leverages the frozen parameters of the backbone model, enabling the caching of activations produced during the forward propagation of the same input sequence, thereby facilitating their reuse across multiple epochs (Cai et al., 2023). As discussed in §4.1, the parallel adapters are a lightweight, separate network that takes the intermediate activations from the backbone transformer as input and generates predictions. During the first epoch, when processing a new input sequence, we cache all the input activations required by the Parallel Adapters that are obtained from the LLM backbone, as illustrated in Figure 5(c), highlighted by the red circle. In subsequent fine-tuning epochs using the same input sequence, we can skip the forward propagation through the LLM backbone entirely, since the required activations have already been cached. The combination of Parallel Adapters and activation caching allows efficient fine-tuning of the LLMs without the need for both forward and backward propagation through the backbone network, thereby (1) significantly accelerating the fine-tuning process and (2) reducing the memory footprint by allowing the release of the memory space occupied by the LLM parameters.
@@
5. Collaborative Edge AI System for Efficient Personal LLMs Fine-Tuning
@@
In PAC, we leverage edge devices in physical proximity and associate them as a resource pool to boost in-situ fine-tuning. Specifically, the fine-tuning procedure comprises two phases: (1) In the initial epoch, the backbone of LLMs, enhanced with Parallel Adapters, undergoes fine-tuning across multiple edge devices through a blend of data and pipeline parallelism (§5.1); (2) In subsequent epochs, the activation cache eliminates the necessity for forward propagation within the backbone, thereby enabling the exclusive fine-tuning of our Parallel Adapters utilizing data parallelism (§5.2).
5.1. Resource-Efficient Collaborative Orchestration for LLMs Fine-Tuning
@@
Observation of Data and Pipeline Parallelism at the Edge. When collaborating on LLM fine-tuning among edge devices, the principle question is which type of parallelism should be used. The most common way to train models in parallel is data parallelism (DP) (Hao and Zhang, 2021). However, DP necessitates that each device maintains a replica of the entire model, a requirement difficult to meet for LLMs with extensive parameter sizes, often surpassing the capacity of a single device. Pipeline parallelism (PP) (Ye et al., 2022) is further proposed to address this problem. In PP, the model is partitioned into multiple consecutive stages and each stage is mapped to a separate device. Consequently, PP enables the training of increasingly large models by deploying more devices. Nonetheless, PP encounters scalability constraints as the addition of edge devices results in more stages. This not only results in a significant presence of pipeline bubbles but also amplifies the impact of inter-stage communication latency, thereby hindering efficiency. The above observation motivates us to employ a hybrid parallelism (HP) architecture that incorporates the best of both DP and PP, so as to achieve superior performance and scalability in resource-constrained edge environments.
@@
Figure 6. An instance of hybrid parallelism in PAC.
Hybrid Parallelism Architecture in PAC. As illustrated in Figure 6(a), PAC first divides an LLM into multiple stages where each contains a stage model composed of a set of consecutive transformer layer. Edge devices are allocated into several device groups, each comprising one or more devices. PAC maps each stage to a group, with the stage model replicated across all devices within that group. Throughout the fine-tuning process, a mini-batch is divided into several micro-batches for concurrent processing to enhance parallelism. If a device cluster hosts multiple devices, micro-batches are further subdivided. Each device is responsible for executing the forward (FP) and backward passes (BP) for its assigned stage model and aggregates gradients across all micro-batches for every mini-batch. Upon completing a mini-batch, gradient synchronization within each device group is achieved through AllReduce. Since the majority of parameters in LLMs are frozen, AllReduce synchronizes only the lightweight parallel adapters, ensuring a swift process. We adopt the one-forward-one-backward ($1$F1111B) micro-batch scheduling (Narayanan et al., 2019) which schedules the BP early to release the activation memory produced by FP for reuse. Figure 6(b) depicts a well-structured hybrid parallelism, encompassing FP, BP, and inter-stage communication.
Profiling. To enable parallelism planning, PAC profiler first fine-tunes the target LLM using calibration datasets to record the runtime profile required for planning. We define $t_{f}^{d,l}(\beta)$ and $t_{b}^{d,l}(\beta)$ as the FP and BP execution times for layer $l$ on device $d$ with batch size of $\beta$, respectively. $u_{d}$ denotes the memory budget of device $d$. The size of output activations, input gradients, and weight parameters in bytes will also be collected to calculate memory footprint.
@@
(2) $W(0\rightarrow y,\mathcal{D}_{n},s)=$ $\min_{0\leqslant q<y}\min_{1\leqslant m<n}\max\{W(0\rightarrow q,\mathcal{D}_{n-m},s-1),$
$T(q+1\rightarrow y,\{d_{n-m}\ldots,d_{n-1}\})\},$
@@
where the first term inside the max is the time of the optimally balanced sub-pipeline between layers 00 to $q$ with $n-m$ devices. The second term represents the time required by the single stage comprising layers $q+1$ to $y$ across $m$ devices. The notation $T(x\rightarrow y,\mathcal{D}_{n})$ denotes the time required for a single stage to execute FP and BP in a data-parallel manner across the device group $\mathcal{D}_{n}$:
@@
(3) $T\left(x\rightarrow y,\mathcal{D}_{n}\right)=\left\{\begin{array}[]{ll}+\infty,\quad\text{if}\ \displaystyle\exists_{d\in\mathcal{D}_{n}}m_{d}>u_{d},\\ \displaystyle\max_{d\in\mathcal{D}_{n}}\displaystyle\sum_{l=x}^{y}\left[t_{f}^{d,l}(\frac{B}{n})+t_{b}^{d,l}(\frac{B}{n})\right],\quad\text{else},\end{array}\right.$
where $M$ is the number of micro-batch and $B$ is the micro-batch size. The peak memory footprint of device $d$, denoted as $m_{d}$ is the sum of the memory usage of the LLM parameters, parameter gradients, and activations. Without out-of-memory (OOM) exceptions, total data-parallel execution time is determined by the slowest device If OOM occurs, the time will be set to positive infinity. During the dynamic programming, we will record pipeline planning configurations, including LLM segmentation points and device groupings.

### v2

@@ III System Overview
PAC+ is a resource efficient collaborative framework for personal LLMs fine-tuning across multiple edge devices. Figure 4 illustrates the PAC+ workflow. PAC+ first constructs a lightweight Parallel Adapter for parameter-efficient LLM fine-tuning, developed as a lightweight version of the target transformer-based LLM. Next, PAC+ optimize resource efficiency during fine-tuning by pre-processing both the LLM and Parallel Adapters. This includes quantizing LLM parameters to a lower bit-width representation for reduced memory footprint, and applying initialization techniques such as structural pruning or knowledge distillation to accelerate convergence (Step 1). PAC+ then integrate the Parallel Adapters module into the LLM, freezing the LLM backbone parameters and fine-tuning only the lightweight adapter for parameter efficient adapting (Step 2). PAC+ profiler fine-tunes the LLM using a calibration dataset on edge devices to record the runtime profile required for parallelism planning (Step 3). PAC+ planner then takes the profiling results as input and generates planning configurations, including LLM partitioning points and device grouping strategies. These configurations comprehensively addresses challenges including memory budget, limited communication capacity, and resource heterogeneity (Step 4). The parallel configurations generated by the PAC+ planner are then applied to the edge devices, enabling time, memory, and parameter-efficient hybrid data and pipeline parallelism fine-tuning of the target LLM (Step 5). Since the LLM backbone parameters remain fixed, the intermediate activations generated by the backbone model are invariant for a given input sequence. The PAC+ maintains a cache of these invariant activations. Through leveraging the cached activations, the efficiency of the fine-tuning process can be accelerated (Step 6).
IV Time, Memory and Parameter Efficient Fine-Tuning Algorithm
IV-A Fine-Tuning LLMs with Parallel Adapters
@@
Observation and Key Insight. As discussed in §II, while techniques such as LoRA [16] and Adapters [13] reduce the number of parameters that need to be updated during fine-tuning, they do not significantly reduce the computational and memory requirements during the training on edge devices. This is because the parameters being updated are still inside the LLM backbone. To calculate the gradients for backpropagation, the full backward passes through the entire pre-trained model are still necessary, as illustrated in Figure 5(a) and (b). In the research field of AI, side-tuning [51] is a specialized fine-tuning technique. It adds a trainable side network that runs in parallel to the backbone model, with the side network’s representation summed with the backbone’s output in the final layer. Crucially, side-tuning only updates the side network, without backpropagating through the backbone model.
@@
Figure 5: Comparison between LLMs fine-tuning with LoRA, Adapters, and our Parallel Adapters.
@@
Parallel Adapters Architecture. In light of side-tuning, we employ a time and memory efficient personal LLMs fine-tuning technique with Parallel Adapters. The overall structure is illustrated in Figure 5(c). Specifically, we decouple conventional Adapters [13] from the LLM backbone, avoiding their integration at the end of each transformer layer. Instead, we provide a dedicated parallel highway for our trainable adapters network, which takes intermediate activations from the backbone transformer as input and generates the final predictions. In this way, backpropagation through the LLM backbone is free, reducing memory demands for massive activations and computational burdens, thereby enhancing time and memory efficiency over techniques like Adapters and LoRA. Observing Figure 5(c), the training paradigm of our Parallel Adapters can be viewed as a shift from fine-tuning the original LLM backbone to fine-tuning a lightweight proxy network that operates in parallel with the backbone. In our design, this lightweight proxy network (i.e., Parallel Adapter) is developed as a lightweight version of the backbone transformer-based LLM [10, 37]. Additionally, our adapter module also demonstrates broad compatibility with established LLM fine-tuning architectures, including the commonly used linear layers for upward and downward projections [13].
Figure 6 provides a detailed illustration of the workflow of our Parallel Adapters architecture. To ensure the lightweight and resource-efficient nature of our proxy network, the hidden dimension of our Parallel Adapters will be $\frac{1}{r}$ times of the original hidden dimension in backbone. Consider an LLM backbone with $L$ layers, where the intermediate outputs $\mathrm{b}_{1},\mathrm{b}_{2},\ldots,\mathrm{b}_{L}$ each contain $n$ tokens, with a hidden dimensionality of $d$, such that $\mathbf{b}_{i}\in\mathbb{R}^{n\times d}$. We denote the embedding input sequence as $\mathrm{b}_{0}\in\mathbb{R}^{n\times d}$. Assuming adapters are inserted after every transformer layer of backbone LLM, Parallel Adapters consist of $L$ adapters, yielding $L$ intermediate outputs $\mathrm{a}_{1},\mathrm{a}_{2},\ldots\mathrm{a}_{L}$, $\mathbf{a}_{i}\in\mathbb{R}^{n\times\frac{d}{r}}$. To align the dimension of $\mathrm{b}_{i+1}$ with $\mathrm{a}_{i}$, we learn a linear projection $\mathbf{W}_{down}^{i}\in\mathbb{R}^{d}\to\mathbb{R}^{\frac{d}{r}}$ to downsample ($\times\frac{1}{r}$) the intermediate activations from the LLM backbone, resulting in $\frac{\mathrm{b}_{i+1}}{r}\in\mathbb{R}^{n\times\frac{d}{r}}$. Next, we combine the downsampled LLM activations $\frac{\mathrm{b}_{i+1}}{r}$ with the intermediate results of the Parallel Adapters $\mathrm{a}_{i}$ using a learnable parameter $\lambda_{i}$: $\lambda_{i}\frac{b_{i+1}}{r}+(1-\lambda_{i})a_{i}$. This combined result is then used as the input to the $i$-th adapter, with $\lambda_{i}$ initialized to 0.5. At the end of the Parallel Adapters, we learn a linear projection $\mathbf{W}_{up}\in\mathbb{R}^{\frac{d}{r}}\to\mathbb{R}^{d}$ to upsample the final outputs to match the dimension of the original language model head. Our evaluation in §VI reveals that Parallel Adapters can achieve comparable model performance to mainstream fine-tuning techniques while being more resource-efficient and better suited for resource-constrained edge environments.
Figure 6: Detailed illustration of the Parallel Adapters architecture.
IV-B Activation Cache for Parallel Adapters
@@
Fine-Tuning Parallel Adapters with PAC+ Activation Cache. Our key idea leverages the frozen parameters of the backbone model, enabling the caching of activations produced during the forward propagation of the same input sequence, thereby facilitating their reuse across multiple epochs [2]. As discussed in §IV-A, the parallel adapters are a lightweight, separate network that takes the intermediate activations from the backbone transformer as input and generates predictions. During the first epoch, when processing a new input sequence, we cache all the input activations required by the Parallel Adapters that are obtained from the LLM backbone, as illustrated in Figure 5(c), highlighted by the red circle. In subsequent fine-tuning epochs using the same input sequence, we can skip the forward propagation through the LLM backbone entirely, since the required activations have already been cached. The combination of Parallel Adapters and activation caching allows efficient fine-tuning of the LLMs without the need for both forward and backward propagation through the backbone network, thereby (1) significantly accelerating the fine-tuning process and (2) reducing the memory footprint by allowing the release of the memory space occupied by the LLM parameters.
Figure 7: Illustration of weight initialization for Parallel Adapters.
IV-C Analysis of Weight Initialization for Parallel Adapters
Observation and Insights. As outlined in §IV-A, our Parallel Adapters fine-tune a lightweight proxy network alongside the original LLM backbone. Proper initialization of the proxy network’s weights is essential for accelerating convergence and reducing computational cost and latency, particularly in resource-constrained edge environments.
We begin by observing the PEFT technique employed by LoRA [16]. LoRA introduces low-rank matrices $A$ and $B$, representing the weight update of the original model as the product of these two matrices, i.e., $\Delta W=BA$. In forward propagation, the model output is expressed as $h=Wx+\Delta Wx=Wx+BAx$. In backward propagation, the LLM backbone weights $W$ are frozen, while the parameters of the low-rank matrices $A$ and $B$ are updated. At the beginning of fine-tuning, LoRA initialized the learnable parameter $\Delta W=0$, with matrix $A$ initialized using a random Gaussian distribution with zero mean, and matrix $B$ initialized as a zero matrix. The rationale behind this initialization strategy is to ensure that the PEFT initial state remains as closely aligned as possible with the pre-trained model, minimizing random perturbations to the model output during the early stages of training [13, 16]. This helps ensure a smooth transition from the pre-trained state to the fine-tuning phase.
We design our initialization scheme based on this insight. As shown in Figure 6, directly applying commonly used random Gaussian initialization or zero initialization to our lightweight parallel adapters results in a significant discrepancy between the initial outputs of the proxy network and the original pre-trained LLM, causing substantial perturbations during the early stages of model fine-tuning. Therefore, we aim for the initialization technique we employ to align the initial states of the lightweight proxy network with frozen LLM backbone. Given that our proxy network is designed as a lightweight version of the backbone LLM, its initialization weights can be derived from model compression techniques, specifically structural pruning and knowledge distillation. In this work, we design and analyze two initialization methods based on above two techniques.
Structural Pruning-Based Initialization. Structural pruning refers to the process of simplifying architectural components, such as dense layers, attention heads, or convolutional channels, while preserving their overall functionality [23, 28, 6]. It is widely employed to achieve model compression, effectively reducing memory footprint and accelerating inference. We observe that the concept of structural pruning can also be applied to initialize the parameters of our Parallel Adapters, where the pruning algorithm takes the backbone LLM as input and outputs its lightweight version for initialization. We implemented the pruning algorithm in our study using Torch-Pruning [6, 38], a popular open-source toolkit for structural pruning atop PyTorch. The fundamental insight of the Torch-Pruning toolkit [38] lies in using a practical norm-based criterion to quantify parameter importance and prune low-significance weights.
Knowledge Distillation-Based Initialization. Knowledge distillation is another widely used model compression technique, where knowledge is transferred from larger, more competent teacher models to smaller student models that are more suitable for resource-efficient deployment [12, 8, 1]. Similarly, in the context of our task, the backbone LLM can be regarded as the larger teacher model, while the Parallel Adapters function as the smaller student model. Specifically, we first prompt the teacher model to generate labels along with rationales justifying the labels for an open dataset. We then leverage these labels, in addition to the rationales, to train smaller student models [15]. We implemented our distillation method based on a popular distillation toolkit open-sourced by Google Research [14]. In practice, since the distillation process does not require the use of private user data, it can be conducted in cloud data centers equipped with more powerful computational resources.
Figure 8: The workflow of mixed-precision computation performed by our Parallel Adapters.
IV-D Further Memory Efficiency through Fine-Tuning with Quantized LLMs
The memory footprint of LLM fine-tuning is primarily contributed by the backbone LLM parameters, intermediate activations, and parameter gradients. Our Parallel Adapters design reduces gradient memory overhead by significantly lowering the number of trainable parameters and minimizes intermediate activation memory overhead by avoiding backpropagation through the backbone model. However, Parallel Adapters do not account for the memory footprint of these substantial backbone model parameters, which can be substantial, especially for LLMs with parameter counts often exceeding hundreds of millions. Quantization reduces memory usage by mapping model weights from higher (e.g., 32-bit or 16-bit floating points) to lower (e.g., 8-bit or 4-bit integers) bit-width representations. This technique is widely adopted for model compression in LLMs serving and fine-tuning [3, 4, 5, 52]. For example, quantizing a 32-bit floating-point (FP32) tensor $\mathbf{X}^{32bit}$ into an 8-bit integer (INT8) tensor $\mathbf{X}^{8bit}$:
$\mathbf{X}^{8bit}=$ $\mathrm{round}\left(\frac{127}{\mathrm{absmax}(\mathbf{X}^{32bit})}\mathbf{X}^{32bit}\right)$ (1)
$=$ $\mathrm{round}\left(c^{32bit}\cdot\mathbf{X}^{32bit}\right),$
where $127$ is the maximum value of the INT8 data type and $\mathrm{absmax}(\mathbf{X}^{32bit})$ represents the maximum absolute value in $\mathbf{X}^{32bit}$. $c^{32bit}$ is the quantization constant for 32-bit tensor $\mathbf{X}^{32bit}$. Correspondingly, dequantization is given by:
$\mathrm{dequant}(c^{32bit},\textbf{X}^{8bit})=\frac{\textbf{X}^{8bit}}{c^{32bit}}=\textbf{X}^{32bit}.$ (2)
This quantization method can be intuitively understood as mapping the range of values representable by FP32 to the range of INT8. However, the issue with this approach is that the presence of large-magnitude values (i.e., outliers) in the input tensor leads to inefficient utilization of the available bit combinations, resulting in precision loss. To address the outlier issue, recent pioneering works [4] inspire the adoption of block-wise quantization algorithms to reduce quantization loss from large-magnitude outliers in the input tensor. We divide the input tensor $\mathbf{X}$ into contiguous blocks and quantize them independently with Equation 1.
Figure 9: Different parallelism plans of a computation graph, where different colors represent different devices and dashed boxes represent stages. For instance, data parallelism involves ingesting data batches into four devices.
In our work, we introduce quantization techniques [4] to reduce the significant memory footprint of backbone LLM parameters and perform our fine-tuning algorithm in a mixed-precision computation paradigm, as illustrated in Figure 8. We have one storage data type (typically a lower-bit data type, e.g., INT8) and a computation data type (typically a higher-bit data type, e.g., FP32). Initially, we quantize the backbone LLM to an 8-bit format for memory-efficient storage, while retaining the lightweight Parallel Adapters in 32-bit precision. During the forward and backward passes, we dequantize the backbone LLM from storage data type to the computation data type to perform FP32-based tensor operations. This mixed-precision design simultaneously offers substantial memory reduction (INT8) on the backbone LLM while provisioning high-precision (FP32) tensors for fine-tuning operations.
@@ IV-D Further Memory Efficiency through Fine-Tuning with Quantized LLMs
V Collaborative Edge AI System for Efficient Personal LLMs Fine-Tuning
@@
In PAC+, we leverage edge devices in physical proximity and associate them as a resource pool to boost in-situ fine-tuning. Specifically, the fine-tuning procedure comprises two phases: (1) In the initial epoch, the backbone of LLMs, enhanced with Parallel Adapters, undergoes fine-tuning across multiple edge devices through a blend of data and pipeline parallelism (§V-A); (2) In subsequent epochs, the activation cache eliminates the necessity for forward propagation within the backbone, thereby enabling the exclusive fine-tuning of our Parallel Adapters utilizing data parallelism (§V-B).
V-A Resource-Efficient Collaborative Orchestration for LLMs Fine-Tuning
@@
Observation of Data and Pipeline Parallelism at the Edge. When collaborating on LLM fine-tuning among edge devices, the principle question is which type of parallelism should be used. We analyze different parallelism plans in Figure 9. The most common way to train models in parallel is data parallelism (DP) [11]. However, DP necessitates that each device maintains a replica of the entire model, a requirement difficult to meet for LLMs with extensive parameter sizes, often surpassing the capacity of a single device. Pipeline parallelism (PP) [47] is further proposed to address this problem. In PP, the model is partitioned into multiple consecutive stages and each stage is mapped to a separate device. Consequently, PP enables the training of increasingly large models by deploying more devices. Nonetheless, PP encounters scalability constraints as the addition of edge devices results in more stages. This not only results in a significant presence of pipeline bubbles but also amplifies the impact of inter-stage communication latency, thereby hindering efficiency. The above observation motivates us to employ a hybrid parallelism (HP) architecture that incorporates the best of both DP and PP, so as to achieve superior performance and scalability in resource-constrained edge environments.
@@
Figure 10: An instance of hybrid parallelism in PAC+.
Hybrid Parallelism Architecture in PAC+. As illustrated in Figure 10(a), PAC+ first divides an LLM into multiple stages where each contains a stage model composed of a set of consecutive transformer layer. Edge devices are allocated into several device groups, each comprising one or more devices. PAC+ maps each stage to a group, with the stage model replicated across all devices within that group. Throughout the fine-tuning process, a mini-batch is divided into several micro-batches for concurrent processing to enhance parallelism. If a device cluster hosts multiple devices, micro-batches are further subdivided. Each device is responsible for executing the forward (FP) and backward passes (BP) for its assigned stage model and aggregates gradients across all micro-batches for every mini-batch. Upon completing a mini-batch, gradient synchronization within each device group is achieved through AllReduce. Since the majority of parameters in LLMs are frozen, AllReduce synchronizes only the lightweight parallel adapters, ensuring a swift process. We adopt the one-forward-one-backward (11F11B) micro-batch scheduling [30] which schedules the BP early to release the activation memory produced by FP for reuse. Figure 10(b) depicts a well-structured hybrid parallelism, encompassing FP, BP, and inter-stage communication.
Profiling. To enable parallelism planning, PAC+ profiler first fine-tunes the target LLM using calibration datasets to record the runtime profile required for planning. We define $t_{f}^{d,l}(\beta)$ and $t_{b}^{d,l}(\beta)$ as the FP and BP execution times for layer $l$ on device $d$ with batch size of $\beta$, respectively. $u_{d}$ denotes the memory budget of device $d$. The size of output activations, input gradients, and weight parameters in bytes will also be collected to calculate memory footprint.
TABLE II: Table of Notations for Parallelism Planning
Notation Definition
$L$ Total number of layers in the DNN model.
$\mathcal{D}$ An ordered set of all devices involved in training.
$\mathcal{D}_{n}$ The subset of first $n$ devices in $\mathcal{D}$.
$M$ Number of micro-batches in a mini-batch.
$B$ Size of each micro-batch.
$u_{d}$ Memory budget available on device $d$.
$m_{d}$ The sum of the memory footprint of LLM parameters, parameter gradients, and activations of device $d$.
$G_{n}$ A device group involving $n$ edge devices.
$t_{f}^{d,l}(\beta)$ / $t_{b}^{d,l}(\beta)$ Time to perform FP / BP for layer $l$ on device $d$ with a batch size of $\beta$.
$c_{f}^{s}(i)$ / $c_{b}^{s}(i)$ FP / BP communication time between stage $i$ and $i+1$.
$W_{s}$ Balanced paratition configurations for pipeline with $s$ stages.
$e_{f}^{s}(i)$ / $e_{b}^{s}(i)$ Time taken for FP / BP in step $i$ with $W_{s}$.
$L_{b}^{s}$ / $L_{e}^{s}$ / $L_{n}^{s}$ Execution time for the Beginning / Execution / Ending phases in $W_{s}$.
$\text{AR}^{s}(i)$ AllReduce time of stage $i$ in $W_{s}$.
@@
$W(0\rightarrow y,\mathcal{D}_{n},s)=\min_{0\leqslant q<y}\min_{1\leqslant m<n}\max\{W(0\rightarrow q,$ (3)
$\mathcal{D}_{n-m},s-1),T(q+1\rightarrow y,\{d_{n-m}\ldots,d_{n-1}\})\},$
@@
where the first term inside the max is the time of the optimally balanced sub-pipeline between layers 0 to $q$ with nmn-m devices. The second term represents the time required by the single stage comprising layers $q+1$ to $y$ across $m$ devices. The notation $T(x\rightarrow y,\mathcal{G})$ denotes the time required for a single stage to execute FP and BP in a data-parallel manner across the device group $\mathcal{G}$. In allocating samples of a mini-batch to devices within $\mathcal{G}$, the goal is to minimize the inference latency determined by the slowest device while ensuring that individual devices do not encounter out-of-memory (OOM) exceptions. We define $H_{x\rightarrow y}(b,\mathcal{G}_{n})$ as the optimal time required by the slowest device in the device group $\mathcal{G}_{n}=\{d_{0},d_{1},...,d_{n-1}\}$ to execute the stage model (layers from $x$ to $y$) when distributing $b$ samples across the group. To solve the aforementioned sample dispatching problem, we also employ a dynamic programming algorithm to search for the optimal sample dispatching strategy. The transition equation for the dynamic programming algorithm can be written as:
@@
$H_{x\rightarrow y}\left(b,\mathcal{G}_{n}\right)=\min_{0\leq i\leq b}\max\left\{\begin{array}[]{ll}H_{x\rightarrow y}(b-i,\mathcal{G}_{n}\setminus\{d_{n-1}\}),\\ \displaystyle\sum_{l=x}^{y}\left[t_{f}^{d_{n-1},l}(i)+t_{b}^{d_{n-1},l}(i)\right].\end{array}\right$. (4)
By solving Equation 4 with dynamic programming, we ultimately obtain $H_{x\rightarrow y}(B,\mathcal{G})$, which represents the desired value of $T(x\rightarrow y,\mathcal{G})$, where $B$ denotes the micro-batch size. We define the peak memory footprint of device $d$, denoted as $m_{d}$, as the sum of memory used by the LLM parameters, parameter gradients, and intermediate activations. Increasing the number of samples allocated to a device enlarges the intermediate activations, potentially leading to OOM exceptions. To exclude OOM cases during the allocation algorithm, the FP and BP time for device $d$ is set to positive infinity. If the algorithm ultimately yields a result of positive infinity for $H_{x\rightarrow y}(B,\mathcal{G})$, it indicates that the collective memory of this group of edge devices cannot accommodate the target stage model. During the dynamic programming, we will record pipeline planning configurations, including LLM segmentation points, micro-batch splitting results, and device groupings strategies.
Input: The target LLM model with $L$ layers. An ordered set of all devices $\mathcal{D}$.
Output: The optimal planning configurations $W_{\sigma}$.
1
2for $s\leftarrow 1$ to $min(L,|\mathcal{D}|)$ do
3  for $y\leftarrow 1$ to $L$ do
4     for $n\leftarrow 1$ to $|\mathcal{D}|$ do
5        for $q\leftarrow 0$ to $y$ do
6           for $m\leftarrow 0$ to $n$ do
7              Get $T(q+1\rightarrow y,\{d_{n-m},...,d_{n-1}\})$ by solving Eq. (4);
8
9           end for
10
11        end for
12       Update $W(0\rightarrow y,\mathcal{D}_{n},s)$ with Eq. (3);
13
14     end for
15
16  end for
17 Calculate $L_{b}^{s}$, $L_{e}^{s}$, $L_{n}^{s}$ using Eq. (5) and (6);
18  Update the optimal stage number $\sigma$ with Eq. (7);
19
20 end for
21Return $W_{\sigma}$;
Algorithm 1 Planning for Hybrid Parallelism